OpenStudy (bianeymariah):

Select the correct inequality to make a true statement? 2/3 ? 3/6 ?={=} {>}, {<}

OpenStudy (3mar):

\[\frac{ 2 }{ 3 }>\frac{ 3 }{ 6 }\]

OpenStudy (3mar):

This is because \[\frac{ 2 }{ 3 }=\frac{ 4 }{ 6 }\]

OpenStudy (3mar):

Thanks for the medal! You are welcome any time

OpenStudy (openstudygirl2):

easy 4/6 the answer did you not know that